OpenStudy (saifoo.khan):

@mominanisar If m/n = .75, then what is 3m + 2n?

OpenStudy (saifoo.khan):

m/n = 0.75 = 75/100 = 3/4 m : n 3 : 4

OpenStudy (saifoo.khan):

Now try it out..

OpenStudy (anonymous):

k thanks.

OpenStudy (saifoo.khan):

Np.